### v1

# Sliding Speed Influences Electrovibration-Induced Finger Friction Dynamics on Touchscreens

Abstract
Electrovibration technology can render tactile textures on capacitive touchscreens by modulating friction between the finger and the screen through electrostatic attraction force generated by applying an alternating voltage signal to the screen. This signal should be carefully calibrated for realistic and robust texture rendering. However, this process is challenging due to variations in sliding speed, applied force, and individual skin mechanics, which affect friction in complex and unpredictable ways. Here, we investigate how exploration conditions affect electrovibration-induced finger friction on touchscreens and the role of skin mechanics in this process. Ten participants slid their index fingers across an electrovibration-enabled touchscreen at five sliding speeds ($20\sim 100$ mm/s) and applied force levels ($0.2\sim 0.6$ N) while we measured contact forces and skin accelerations. The touchscreen was excited with amplitude-modulated voltage signals across frequencies relevant to touch. We modeled the finger-touchscreen friction response as a first-order system and the skin mechanics as a mass-spring-damper system. Our results showed that the sliding speed influenced the cutoff frequency of the friction response as well as the moving mass and stiffness of the finger for the tested exploration ranges. Specifically, for every 1 mm/s increase in speed, the cutoff frequency, the finger moving mass, and stiffness increased by 13.8 Hz, $3.23\times 10^{-5}$ kg, and 4.04 N/m, respectively. Further correlation analysis revealed that finger stiffness affected the cutoff frequency more than the moving mass. Finally, we developed a practical model for electrovibration-induced finger friction on touchscreens that accounts for sliding speed variations, paving the way for delivering consistent haptic feedback through electrovibration.
keywords:
Surface haptics , Electrovibration , Skin mechanics , System identification , Texture rendering
journal: arXiv
[…]
[label1]organization=Dept. of Cognitive Robotics, Mechanical Engineering, TU Delft,addressline=Mekelweg 5, city= Delft, postcode=2628 CD, country=The Netherlands
[…]
[label2]organization=Dept. of Control & Operations, Aerospace Engineering, TU Delft,addressline=Kluyverweg 1, city= Delft, postcode=2629 HS, country=The Netherlands
{graphicalabstract}[Uncaptioned image]
{highlights}
Sliding speed impacts electrovibration-induced finger friction bandwidth
Applied force does not affect electrovibration-induced finger friction dynamics
Skin stiffness influences finger-electrovibration bandwidth more than mass does
A model was proposed to compensate for speed-based variations in electrovibration friction
1 Introduction
Imagine a touchscreen that does not just sense your touch but lets you feel the textures – whether it is the roughness of tree bark or the smoothness of silk – beneath your fingers. The field of surface haptics aims to generate artificial tactile sensations, such as mimicking the touch experience of interacting with physical textures on flat surfaces, by modulating friction or vibrations between the user’s finger and a flat surface [1]. Recreating the experience of touching physical surfaces – known as data-based texture rendering – involves recording the contact forces or skin vibrations when a finger interacts with physical textures. Then, these finger-surface interaction recordings serve as input for surface haptic devices equipped with actuators, which convert electrical signals into friction or vibration outputs.
Electrovibration is one such technology that can modulate friction between a touchscreen and a user’s finger via electrostatic forces [2]. When a finger interacts with the touchscreen, it functions as a parallel plate capacitor, where the screen’s non-conductive coating, the outermost layer of the skin (stratum corneum), and an air gap between the skin and the touchscreen (caused by the finger ridges and touchscreen’s rough surface) act as the dielectric layers [3, 4, 5]. The touchscreen’s conductive layer beneath the non-conductive coating and the skin’s inner tissues serve as two parallel conductors. Applying an alternating high-voltage signal to the conductive layer of the touchscreen generates an electrostatic force between the finger and the touchscreen, pulling the skin toward the surface in response to the input voltage signal. When the finger moves, this attraction and release of the skin alter the friction between the finger and the screen, creating the sensation of texture on an otherwise smooth glass surface. By using recorded contact forces from real-world finger-surface interactions as input signals, electrovibration displays can replicate virtual texture sensations that closely match their physical counterparts [6, 7]. Electrovibration outperforms other surface haptic technologies for its texture rendering capability, as it has a wide bandwidth, generates localized skin vibrations across the touchscreen, and the produced sensations are independent of the screen size [8].
Despite its advantages, rendering realistic textures with electrovibration remains challenging. Electrostatic force generation induces a mismatch between the frequency spectra of the input voltage and the resulting friction signals [9, 10]. One reason for this mismatch is the nonlinear behavior of the electrostatic forces in a parallel plate capacitor system, where the generated force is proportional to the square of the input voltage (see Section 2 for more details). Additionally, the finger-surface contact acts as an electrical high-pass filter, attenuating low-frequency components in the generated forces [11, 12, 3, 13, 14]. Moreover, the mechanical behavior of the fingertip skin, which was shown to behave as a low-pass filter [15, 16], further contributes to this mismatch.
Achieving accurate texture rendering with electrovibration devices requires calibrating the input voltage to account for the input-output spectral discrepancy. Meyer et al. [17] addressed the electrical high-pass behavior by modulating the amplitude of a high-frequency voltage signal with the desired frequency spectrum. Later, Shultz et al. [18] solved the input squaring by applying the square root operation to the envelope of the amplitude-modulated signal. Subsequently, Grigori et al. [19] built a closed-loop controller, dynamically measuring the friction force and controlling input current to maintain a constant voltage, with amplitude modulation, and achieved a theoretical 90% texture tracking by the controller. However, despite the large variances observed in the experiments, the controller was designed with a single gain term for the input current-to-friction relationship. These variations were attributed to differences in exploration conditions and skin mechanics; failure to account for these could make the controller unstable. Since addressing the input current-to-friction variation was complex, the same authors in the later study [7] used an open-loop system and constant exploration conditions for texture rendering and achieved 71% perceptual texture similarity.
Although often overlooked, sliding speed and applied force greatly contribute to the spectrum mismatch for the desired friction on electrovibration devices [20]. The soft multilayered structure of the fingertip [21] influences its dynamic response [22], with behavior highly dependent on both speed and force [15, 23]. Previous studies have demonstrated that the frequency composition of skin vibrations shifts toward higher frequencies with increasing speed [24, 25]. Similarly, the magnitude of the skin vibrations reduces with an increase in finger normal force [26]. Additionally, the skin’s mechanical properties – mass, damping, and stiffness – rise with the increase in applied force, affecting its vibration spectrum [16]. Applied normal force and sliding speed also cause variations in the air gap, leading to variations in generated electrostatic forces [27, 28, 9, 13]. However, despite these observations, a thorough analysis of how exploratory forces and speeds affect electrostatic forces and how to compensate for these effects in texture rendering has yet to be performed.
In this paper, we investigate how exploration conditions influence the friction between a finger and an electrovibration display, and the role that skin mechanics play in this interaction. We measured the generated friction forces and skin vibrations when 10 participants explored an electrovibration display under controlled conditions, varying both applied force (0.2 N, 0.3 N, 0.4 N, 0.5 N, and 0.6 N) and sliding speed (20 mm/s, 40 mm/s, 60 mm/s, 80 mm/s, and 100 mm/s). The display was driven with 15 logarithmically spaced input frequencies (30 Hz $\sim$ 2 kHz) modulated using the technique described by Shultz et al. [18]. We then identified and modeled the electromechanical behavior between input voltage and measured friction force to analyze how the model parameters change as a function of applied forces and speeds. Additionally, since skin biomechanical properties vary among individuals, we analyzed the model’s variability across 10 participants. Finally, we investigated the relationship between the skin’s mechanical properties and the corresponding electro-mechanical response. We believe that this study offers valuable new perspectives for developing more realistic texture interactions on touchscreens.
2 Methodology
Figure 1: An illustration of how a message signal, $V_{m}(t)$ is modulated with a carrier signal, $C(t)$, to generate the input voltage, $V_{i}(t)$. The input signal is applied to a touchscreen that generates electrovibration force, $F_{e}(t)$, resulting in friction, $F_{f}(t)$, and skin vibration as an acceleration signal, $a(t)$, during finger sliding.
2.1 Electrovibration
Electrovibration modulates friction between a capacitive touchscreen and a moving finger using electrostatic forces. The resulting friction force, $F_{f}$, is influenced by both the applied normal force exerted by the finger, $F_{n}$, and the electrostatic force, $F_{e}$. The relation $F_{f}=\mu(F_{n}+F_{e})$ describes the friction force, where $\mu$ represents the friction coefficient. The $F_{e}$ depends on the applied input voltage, $V_{i}$, and is given by the equation [3]:
$F_{e}=\frac{A\epsilon_{o}\epsilon_{g}}{2}\left(\frac{V_{i}}{d_{g}}*\left|\frac{Z_{g}(\omega)}{Z_{g}(\omega)+Z_{d}(\omega)+Z_{sc}(\omega)}\right|\right)^{2}$ (1)
where $A$ is the finger’s contact area, $\epsilon_{o}$ is the permittivity of free space, $\epsilon_{g}$ and $d_{g}$ are the permittivity and thickness of the air gap. $Z_{g}(\omega)$, $Z_{d}(\omega)$, and $Z_{sc}(\omega)$ represent the electrical impedance of the air gap, the dielectric layer of the touchscreen, and the stratum corneum, respectively.
According to Equation (1), the electrovibration force is proportional to the square of the input voltage, expressed as $F_{e}=kV_{i}^{2}$, where $k$ is a constant [13]. This quadratic relationship introduces nonlinear behavior, causing the output force to oscillate at twice the input frequency. This frequency doubling expands the input spectrum, causing a discrepancy between the input voltage signal and the output skin vibration. See A for derivation.
Another challenge in electrovibration is charge leakage from the conductor to the finger, caused by the electrical properties of the finger-surface contact interface [12, 29, 14]. When a low-frequency voltage input is used as an excitation signal, charges have enough transit time to move from the touchscreen’s conductor to the finger through the ridges and vice versa. This reduces the voltage drop across the air gap, reducing the electrovibration force intensity. This leakage significantly attenuates low-frequency components, causing the input voltage to output friction response to exhibit a high-pass filtering effect [11, 9].
As mentioned in Section 1, eliminating this nonlinearity and any charge leakage is necessary for minimizing the discrepancy between the input voltage and output friction for realistic texture rendering. Due to these physical phenomena, the input voltage spectrum is attenuated at low frequencies and spreads, causing the output friction spectrum to mismatch with the input voltage spectrum. As humans are sensitive to tactile frequencies up to 1 kHz [30], any output friction distortions below 1 kHz may cause a different texture feel than intended, reducing the virtual texture realism.
One way to mitigate charge leakage and nonlinearity is through amplitude modulation followed by a square root operation on its envelope. In amplitude modulation, the input message signal, $V_{m}(t)$, is modulated with a high-frequency carrier, $C(t)=\cos(2\omega_{c}t)$[31]. The bandwidth of the message signal is from 1 Hz to 1 kHz, $\omega_{m}=2\pi f_{m}\leq 1$ kHz, where $f_{m}$ and $\omega_{m}$ are the message signal’s frequency and angular frequency. The carrier frequency is typically above 7 kHz so that humans perceive the envelope of the signal and not the carrier itself, $\omega_{c}=2\pi f_{c}\geq 7$ kHz, where $f_{c}$ and $\omega_{c}$ are the carrier signal’s frequency and angular frequency. Also, the carrier distorts the message signal if it is too close to the message signal’s bandwidth. The amplitude-modulated input signal, $V_{i}(t)$, has three frequency components – $\omega_{c}-\omega_{m}$, $\omega_{c}$, and $\omega_{c}+\omega_{m}$ – in the pass-band of the high-pass input voltage to output friction response. During finger electrovibration interaction, the doubling of the input frequency components results in output frequency components at $2\omega_{m}$ and $2\omega_{c}$. The finger attenuates any frequency component above 1 kHz [32], resulting in output friction forces at $2\omega_{m}$. To ensure the output friction occurs at $\omega_{m}$, we shift the message by its minimum, $|V_{m}(t)|$, and apply a square root operation to the envelope of amplitude-modulated signal as shown in Equation (2). Refer to A for further details and mathematical derivations.
$V_{i}(t)=\sqrt{V_{m}(t)+|\min(V_{m}(t)|}\cos(2\omega_{c}t)$ (2)
An illustration describing the process of how an amplitude-modulated input voltage applied to a touchscreen generates electrovibration forces, resulting in friction and skin vibration, is shown in Figure 1.
2.2 Experimental setup
Figure 2: Experimental setup: (1) grounding strap, (2) force sensor and touchscreen, (3) accelerometer, (4) data acquisition card connector, (5) amplifier, (6) monitor displaying the experiment graphical user interface.
We used a custom setup (see Figure 2). The setup featured a touchscreen (SCT3250, 3M), which we cut to a size of $35\times 70$ mm. The electrostatic forces were generated by sending voltage signals to the conductive layer of this touchscreen. The voltage signals were designed using a PC (Precision T5820, Dell) and generated by a data acquisition card (PCIe-6323, NI) with a sampling frequency of 20 kHz. These signals were then amplified through a high-voltage amplifier (9200A, Tabor Inc.) with a gain of 50 before sending them to the touchscreen.
The touchscreen was placed on a force sensor (Nano43 Titanium, ATI) using double-sided tape (Tesa PRO) to measure the generated forces in three dimensions. We mounted the force sensor on a large aluminum block with screws. This entire assembly was placed on an aluminum optical breadboard table (MB60120/M, Thorlabs), supported by a sturdy frame (PFM52502, Thorlabs). A monitor (UltraSharp 24, Dell) displayed graphical information to the participants. The data acquisition card collected data from the force sensor at a sampling rate of 20 kHz. A 2D infrared position sensor (NNAMC1580PCEV, Neonode) tracked the finger’s position with a resolution of 0.1 mm and a sampling rate of 60 Hz. The sensor was mounted on a 3D-printed support, ensuring its active plane was positioned above the touchscreen. We enclosed the touchscreen in a box with a 70$\times$80 mm cutout for finger exploration. We also measured the three-dimensional skin vibrations via an accelerometer (ADXL358, Analog Devices) attached to the participants’ fingernails via an accelerometer wax (Model 080A109, PCB Piezotronics). The accelerometer data was also collected through the same data acquisition card with a sampling rate of 20 kHz.
We designed the experimental setup to ensure the mechanical resonance frequency outside the tactile range (1 Hz to 1 kHz). As the experimental setup has peak sensitivity at its resonance frequency, any input force at this frequency would be measured as higher than the original signal intensity. Hence, a resonance frequency below 1 kHz would lead to an error in the measured force signals in the tactile range. To verify the setup’s bandwidth, we struck the touchscreen with an impact hammer (086E80, PCB Piezotronics) along the normal and lateral directions and measured the response using the force sensor beneath the screen. The results indicated a resonance frequency of 1.4 kHz in the normal direction (see Figure 9) and two resonances at 866 Hz and 1740 Hz in the lateral direction (see Figure 10).
2.3 Procedure
We conducted the experiments following the ethical principles of the Declaration of Helsinki. We obtained approval from the TU Delft Human Research Ethics Committee (application number 3469). The study involved five women and twelve men, with an average age of 26.3 years (SD: $\pm$2.36 years). All participants provided informed consent before the experiments.
The aim of the experiments was to record the finger interaction forces and accelerations while the participants explored the electrostatically actuated touchscreen at a 60 contact angle. The cut-out of the enclosure box around the touchscreen ensured that the finger always maintained a constant contact angle with the touchscreen (see Figure 2). During the experiments, a participant moved the index finger of their dominant hand on the touchscreen horizontally (left to right) at five different sliding speeds (20 mm/s, 40 mm/s, 60 mm/s, 80 mm/s, and 100 mm/s) while applying five different normal forces (0.2 N, 0.3 N, 0.4 N, 0.5 N, and 0.6 N). A custom graphical user interface (GUI) guided the participants by displaying two visual square cues: the top square indicated the reference speed, while the bottom square tracked the user’s finger movement (Figure 2). The tracking square changed color to indicate the applied force: green for the target range, yellow for below the target, and red for above the target. If the participant’s mean, minimum, and maximum exploration force and speed deviated by ±$\pm 10\%$ of the target range for more than one second, then the trial was discarded and repeated.
Each participant explored the touchscreen at different speeds and forces while it was excited with 15 logarithmically spaced sinusoidal signals ranging from 30 Hz to 2 kHz. Each signal was amplitude-modulated with a 7 kHz carrier frequency, as defined in Equation (2). The peak-to-peak amplitude of the final input voltage signal was 150 $V_{pp}$.
Each trial lasted 17 seconds and consisted of two phases. During the first 7 seconds, the participant adjusted their finger exploration to the experimental conditions; however, their interaction data was not recorded during this period. In the remaining 10 seconds, we recorded lateral force and skin vibration data. During this period, the participants swiped the touchscreen from left to right six times. After each trial, the process was repeated for subsequent trials.
This experimental configuration resulted in 375 trials (15 frequencies $\times$ 5 speeds $\times$ 5 forces) for each participant. After every 15 trials, we gave the participants a 10-second break. Each participant spent approximately 6 hours completing data collection. Due to the long experiment duration, the participants performed the experiment in multiple sessions.
Some participants did not have good hand-eye coordination and failed to meet the target range for the experimental conditions. We rejected the data from seven participants who spent more than one hour on a trial set of one sliding speed and five exploration forces. In the end, three women and seven men successfully completed the experiment, whose data was further processed.
Figure 3: An example of experimental data collected during the finger-touchscreen interaction at a speed of 40 mm/s and a normal force of 0.2 N. A 30 Hz message signal was shifted with its minimum and modulated with a 7 kHz carrier signal. The resulting amplitude-modulated signal’s envelope was square-rooted and then amplified to 150 V (peak to peak) to excite the touchscreen. The plots show raw data, with the highlighted sections extracted for analysis. The recorded (a) amplitude-modulated input voltage signal, (b) 1D friction signal obtained through dimension reduction of measured lateral forces, and (c) finger vibration as the integrated velocity signal derived from the raw acceleration data.
2.4 Data processing
During the experiments, we recorded the contact forces, finger accelerations, and the input voltage signal after modulation for analysis, as shown in Figure 3(a). The finger’s active exploration was, in general, not exactly a straight line; this generated lateral forces, which were primarily recorded by the force sensor’s y-axis and partially along the x-axis. To capture the generated friction forces accurately, we transformed the two-dimensional lateral force data into a single dimension in the frequency domain [33], extracting the friction signal as shown in Figure 3(b). Similarly, the accelerometer’s x-axis primarily recorded the lateral finger vibrations along the finger exploration direction. To align the accelerometer axes with the force sensor axes, we multiplied two rotation matrices: one along the x-axis at $155^{\circ}$ and another along the y-axis at $180^{\circ}$, accounting for the inverse orientation of the accelerometer on the y-axis and the finger exploration angle of $60^{\circ}$. After multiplying the raw sensor readings by $9.8m/s^{2}$ for conversion to a true acceleration signal, we integrated the signals in time to obtain the velocity data, as shown in Figure 3(c).
From the input voltage signal (amplitude modulated), the friction force, and the finger vibration signal, we extracted sections to identify mathematical models. We used the friction data as a reference and set a threshold condition to determine when the participants made a finger sweep from left to right. If the friction force was above its mean value, then the participant would sweep from left to right. Next, we found the start and end indices of all the left-to-right sweeps made by the participants. We then averaged a particular sweep’s start and end index to find its middle index. We proceeded to extract samples of 1999 long to the left and 2000 long to the right of the middle index. By using the time stamps of the 4000 samples long data, we extracted the corresponding data sections from the amplitude-modulated signal and the finger vibration signal as shown by the highlighted sections in Figure 3. We selected the sample size of 4000 because it is a multiple of the sampling rate 20 kHz, which ensures all test sinusoidal frequencies appear in the Fourier transform. Additionally, at the maximum sliding speed of 100 mm/s, the recorded data was 7500 samples long, so we used 4000 samples for consistency across all exploration conditions.
After extracting the samples for the input amplitude-modulated voltage signals, $V_{i}$, for different finger sweeps, we demodulated them to get the original message signals back to model the input message signal, $V_{m}$, to output friction, $F_{f}$, response. To do this, we applied the Fourier transform on the input amplitude-modulated signal, $V_{i}$, and then we calculated the magnitude and phase of the upper envelope, $\omega_{c}+\omega_{m}$, of the amplitude-modulated signal. Using the magnitude and phase value, we reconstructed the sinusoidal message signal. We repeated the same process for all the sweep data, frequencies, exploration conditions, and participants.
2.5 Dynamic modeling of electrovibration-induced finger friction on a touchscreen
Figure 4: Bode plot of the electrovibration-induced finger friction on a touchscreen, $H_{f-e}$, for input voltage envelope when the display was explored with a sliding speed of 40 mm/s and an applied force of 0.2 N. Each point in the figure represents the Fourier transform when the output friction force, $F_{f}$, is divided by the input message signal, $V_{m}$, at that frequency. Each color shows the sweep order made by the finger at a given trial. The response up to 750 Hz reflects the electrovibration-induced finger friction, while the mechanical response of the measurement setup dominates frequencies beyond 750 Hz. The black curve represents the model incorporating both the finger and setup response. The red curve represents the first-order electrovibration-induced finger friction response after removing the setup’s behavior.
Our goal was to investigate how exploration speed and force influence the electrovibration-induced finger friction on a touchscreen. To achieve this, we modeled the relationship between the input message signal, $V_{m}$, and output friction, $F_{f}$, while assessing the impact of exploration conditions on the model parameters. Given the nonlinear behavior of the fingertip when interacting with a surface [34], we sought to simplify the electrovibration-induced finger friction model by conditionally representing it as a linear system. To check this, we identified the transfer function, $H_{f-e}(s)$, by dividing the Fourier transform of the output friction, $F_{f}$, by the Fourier transform of the input message signal, $V_{m}$, for a particular exploration force and speed. We repeated this process for every finger sweep made by the participant and every test frequency to construct a Bode plot as shown in Figure 4; the input-output dynamics behave like a first-order low-pass up to 750 Hz and then show two resonance peaks at 860 Hz and 1740 Hz. We verified that the two resonance peaks are due to the experimental setup’s hardware; please refer to B for more details. We used the fmincon algorithm in MATLAB to optimize the transfer equation based on the discrete transfer function points found earlier. Then, we removed the setup’s dynamics as given by Equation (23) from the fitted transfer function model. As a result, we found a first-order response for the input message signal, $V_{m}$, to output friction, $F_{f}$, response, as shown in Equation (3).
$H_{f-e}(s)=\frac{F_{f}(s)}{V_{m}(s)}=\frac{K\omega_{o}}{s+\omega_{o}}$ (3)
Here, $F_{f}(s)$ and $V_{m}(s)$ are the output friction and input message signal in the frequency domain, $K$ is the system’s gain, and $\omega_{o}$ is the cutoff frequency.
This first-order behavior aligns with previous findings by Meyer et al. [17], who demonstrated that the message signal of an amplitude-modulated voltage signal and the output friction, $F_{f}$, in finger-electrovibration interaction followed a first-order low-pass response at an arbitrary exploratory force and speed. They modeled the behavior using the device’s quality factor, $Q$, and carrier frequency, $\omega_{c}$, as $H_{f-e}=\frac{Q}{1+{2Qs}/{\omega_{c}}}$. However, to better represent the input message signal, $V_{m}$, to output friction, $F_{f}$, response variation caused by exploration speed and force, we opted for a generic model given by Equation (3). Figure 4 shows an input message signal, $V_{m}$, to output friction, $F_{f}$, response computed for one of our participants at 40 mm/s exploration speed and 0.2 N normal force. The corresponding estimated model parameters are $K=0.0147$ (N/V), and $\omega_{o}=961$ Hz.
Figure 5: Bode plot of the finger mechanical response, $H_{f}$, for input electrovibration-induced finger friction when the display was explored with a sliding speed of 40 mm/s and an applied force of 0.2 N. Each point in the figure represents the Fourier transform when the output finger vibration velocity is divided by the input finger friction force. Each color shows the sweep order made by the finger at a given trial. The model analysis is limited to 750 Hz due to the resonance of the setup.
2.6 Dynamic modeling of finger skin mechanical behavior
Additionally, we modeled finger vibrations caused by the friction signal to investigate how the finger’s mechanical behavior influences electrovibration-induced finger friction on a touchscreen. Here, we considered the measured electrovibration-induced finger friction force, $F_{f}$, as input and finger vibration velocity, $v$, as output. As the finger exhibits a nonlinear behavior [35, 34, 36], we wanted to know if the input friction force, $F_{f}$, to output finger vibration velocity, $v$, can be represented by a linear system. To do this, we divided the Fourier transform of the finger vibration signal by the Fourier transform of the friction signal to obtain the transfer function, $H_{f}$. We repeated the process for every finger sweep and frequency for a particular exploration condition, similar to finger-electrovibration modeling. After plotting the transfer function points using a Bode plot, see Figure 5, we observed that the system showed a second-order response. Using the fmincon algorithm in Matlab, we modeled the system based on the commonly used mass, spring, and damper system in the literature [15, 16, 7]:
$H_{f}(s)=\frac{v(s)}{F_{f}(s)}=\frac{s}{ms^{2}+bs+k}$ (4)
Here, $v(s)$ and $F_{f}(s)$ are the Fourier transform of the finger vibration as velocity signal and friction generated due to finger electrovibration interaction, $H_{f}(s)$ is the transfer function between them, $m$ is finger-moving mass, $b$ is finger damping, and $k$ is finger stiffness. The moving mass of the finger represents the effective mass of the subcutaneous finger pad, while the damping and stiffness correspond to the viscosity and elasticity of the fingertip pulp [15]. Figure 5 shows an example finger mechanical response obtained when the display was explored with 40 mm/s sliding speed and 0.2 N applied force. The finger mechanical model parameters for the response are $m=0.0015$ Kg, $b=1.3$ Ns/m, $k=444$ N/m. Our findings matched the mechanical response, their derived mass, spring, and damping values reported in the previous studies [15, 16, 7], even though there are other techniques to excite and measure the finger response.
3 Results
3.1 Dynamics of electrovibration-induced finger friction on touchscreens
Figure 6: The first-order model parameters of the transfer function, $H_{f-e}$, between the measured friction and input message signal, $V_{m}$, for the tested exploration speeds and forces across participants: (a) gain, $K$, and (b) cutoff frequency, $\omega_{o}$. The results corresponding to each exploration force and speed are shown in separate color-coded box plots. The center lines show the medians; box limits indicate the 25th and 75th percentiles. The whiskers extend to 1.5 times the interquartile range. The filled black circle ($\bullet$) and the plus (+) symbols indicate the outliers and means, respectively. The asterisk (*) and double asterisk (**) with the connected brackets show the statistically significant pairs at ($p<0.05$) and ($p<0.01$), respectively.
The estimated gains and cutoff frequencies of the first-order models fitted to the measured friction signal and input message signal, $V_{m}$, under the tested exploration conditions are shown in Figures 6(a) and 6(b). We performed statistical analyses to examine how sliding speed and force influence electrovibration-induced finger friction. In our analysis, the applied force, $F_{n}$, and sliding speed, $\nu$, were treated as independent variables (main effects), while the model parameters – the gain $K$ and cutoff frequency $\omega_{o}$ – were the dependent variables. We assessed the normality of all data using the Kolmogorov-Smirnov test, which indicated that both the gains and cutoff frequencies were not normally distributed ($p<0.01$). Hence, we applied a generalized linear mixed model (GLMM) to test the influence of the independent variables. To account for the variability in these parameters introduced by different participants, the participant ID number was included as a random effect. The resulting mixed model with a linear approximation for the gain parameter $K$ is given as:
$K=\alpha_{o}+\alpha_{1}F_{n}+\alpha_{2}\nu+\alpha_{3}F_{n}*\nu+\rho+\epsilon$ (5)
Here, $\alpha_{o}$ is the model’s intercept parameter, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are the fixed coefficients corresponding to the effects of force, speed, and the force-speed interaction, $\rho$ is the coefficient for the random between-participant effect, and $\epsilon$ is the model’s residual. The fixed coefficients show the estimated influence of the independent variable on the dependent variable, with other independent variables being constant. Our analysis revealed that the main effects of applied force ($F(1,178)=0.310$, $p=0.578$) and sliding speed ($F(1,214)=0.715$, $p=0.399$), and their interaction ($F(1,210)=0.892$, $p=0.346$) all did not affect the gain. Moreover, between-participant variation also had no significant effect on the measured $K$ values (Estimate: $2.96\times 10^{-5}$, SE: $1.53\times 10^{-5}$, $Z=1.92$, $p=0.054$).
Similarly, the mixed model with linear approximation used for statistical analysis of the cutoff frequency $\omega_{o}$ data is given as:
$\omega_{o}=\beta_{o}+\beta_{1}F_{n}+\beta_{2}\nu+\beta_{3}F_{n}*\nu+\rho+\epsilon$ (6)
Figure 7: Results showing the variation of finger’s mechanical response, $H_{f}$, with sliding speed and applied force across participants. Box plots of finger’s (a) mass, $m$, (b) damping, $b$, and (c) stiffness, $k$. The results corresponding to each applied force and sliding speed are shown in separate color-coded box plots. The center lines show the medians; box limits indicate the 25th and 75th percentiles. The whiskers extend to 1.5 times the interquartile range. The filled black circle ($\bullet$) and the plus (+) symbols indicate the outliers and means, respectively. The asterisk (*) with the connected brackets shows the statistically significant pairs at ($p<0.05$).
Here, $\beta_{o}$ is the model intercept, $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are the fixed coefficients for force, speed, and the force-speed interaction, $\rho$ is the coefficient for the random effect due to participant ID, and $\epsilon$ is the model’s residual. Unlike the statistical analysis results for $K$, the cutoff frequency was significantly affected by the sliding speed ($F(1,211)=25.068$, $p<0.001$). However, also no significant effects of the applied force ($F(1,176)=1.12$, $p=0.291$), the speed-force interaction ($F(1,207)=0.715$, $p=0.399$), and the between-participant effect (Estimate: $4.41\times 10^{4}$, SE: $2.43\times 10^{4}$, $Z=1.816$, $p=0.069$) were found. Hence, only the finger speed significantly affects the cutoff frequency $\omega_{o}$; the corresponding fixed coefficient, $\beta_{2}$, has a value of 13.811 Hz/mm/s, meaning that for every increase in speed by 1 mm/s, the cutoff frequency increases by 13.811 Hz (with respect to an intercept, $\beta_{o}$, of 385.68 Hz). Next to the GLMM analysis, we also performed a post-hoc Wilcoxon signed-rank test to evaluate if the cutoff frequencies across the participants statistically differ with sliding speeds for different forces; see bars at the top of Figure  6(b) for the significantly different pairs. The post-hoc analysis showed a significant difference in cutoff frequency across sliding speeds, particularly between the lower and higher speed ranges. However, as the applied force increased, the cutoff frequency remained insignificant at higher force levels (60–100 mm/s).
Based on our GLMM analysis results, we propose the following empirical model that can be used to predict the electrovibration-induced finger friction dynamics as a function of the sliding speed, $\nu$:
$H_{f-e}(s,\nu)=\frac{K\omega_{o}}{s+\omega_{o}(\nu)}\in K=0.0123;\omega_{o}=385.68+13.811\nu$ (7)
Here, $K=0.0123$ is the average gain across all sliding speeds, applied forces, and participants measured in the experiment, as our GLMM analysis showed no effects of these factors on this gain. However, as the cutoff frequency was influenced by the sliding speed, we simplified the corresponding linear model as $\omega_{o}(\nu)=385.68+13.811\nu$ to describe the variation in finger-display interaction with sliding speed. Note that this model is only valid for an input voltage amplitude of 150 $V_{pp}$; it does not capture variations with amplitude changes.
3.2 Dynamics of finger’s mechanical behavior
Figure 8: Scatter plot showing the Pearson’s correlation between the finger-display interaction parameters—$K$ and $\omega_{o}$—and finger mechanical response parameters—$m$, $b$, $k$. (a) Gain, $K$ vs. finger moving mass, $m$, (b) cutoff frequency, $\omega_{o}$, vs. finger moving mass, $m$, (c) gain, $K$, vs. finger damping, $b$, (d) cutoff frequency, $\omega_{o}$, vs. finger damping, $b$, (e) gain, $K$, vs. finger stiffness, $k$, (f) cutoff frequency, $\omega_{o}$, vs. finger stiffness, $k$. The blue dots indicate the data points, and the dashed line indicates Pearson’s correlation fit. The correlation coefficient, $r$, and significance value, $p$, are shown in the plot.
The mass, damping, and stiffness of the second-order model fitted to the measured finger vibration velocity for input finger friction under different exploration conditions are shown in Figures 7(a)7(b), and 7(c), respectively. Similarly, a Kolmogorov-Smirnov test indicated that these parameters were not normally distributed ($p<0.01$). We assigned the model parameters -— the finger moving mass $m$, finger damping $b$, and finger stiffness $k$ – as dependent variables, the sliding speed and applied force as the main effects, as well as their interaction, and participant ID as a random effect. The GLMM analysis revealed that the finger moving mass was affected by sliding speed ($F(1,205)=7.193$, $p=0.008$) with a fixed coefficient of $3.5\times 10^{-5}$ Kg/mm/s, but not by the applied force ($F(1,176)=0.125$, $p=0.724$) or their interaction ($F(1,201)=1.216$, $p=0.271$). The estimated coefficient values indicate that the finger moving mass increased by $0.0028$ kg from the lowest to the highest sliding speed. The finger damping was not affected by the applied force ($F(1,179)=3.280$, $p=0.071$), sliding speed ($F(1,224)=2.843$, $p=0.093$), or their interaction ($F(1,221)=1.604$, $p=0.207$). The finger stiffness was affected by sliding speed ($F(1,208)=14.465$, $p<0.001$) with a fixed coefficient of $4.036$ N/m/mm/s, but not by the applied force ($F(1,175)=1.148$, $p=0.285$) or their interaction ($F(1,204)=3.79$, $p=0.053$). Based on the obtained coefficient values, the finger stiffness increased by $322.88$ N/m from the lowest to the highest applied force. Moreover, the obtained moving mass (Estimate: $4.99\times 10^{-6}$, SE: $2.41\times 10^{-6}$, $Z=2.068$, $p=0.039$), damping (Estimate: 0.356, SE: 0.180, $Z=1.979$, $p=0.048$), and stiffness (Estimate: $2.36\times 10^{4}$, SE: $1.16\times 10^{4}$, $Z=2.03$, $p=0.042$) values were significantly affected by the participant ID. We also conducted a post-hoc Wilcoxon signed-rank test to reveal for which tested sliding speeds these values significantly differ (see bars above figures in Figures 7(a)7(b), and 7(c)).
3.3 Influence of skin mechanics on finger-display interaction
As a final analysis step, we performed a Pearson’s correlation analysis between the finger-display interaction parameters (the gain, $K$, and cutoff frequency, $\omega_{o}$) and the finger mechanical response parameters (the moving mass, $m$, damping, $b$, and stiffness, $k$) for all the sliding speeds and force, as shown in Figure 8. The results indicate that gain significantly decreased with an increase in finger moving mass, and increased with an increase in finger damping. Similarly, the cutoff frequency significantly increased with higher finger moving mass and finger stiffness.
4 Discussion
In this study, we investigated how exploration conditions influence electrovibration-induced finger friction on touchscreens and the role of skin mechanics in this relation. We measured the contact forces and skin accelerations generated as ten participants moved their index fingers on an electrovibration display under varying exploratory forces and speeds. Using this data, we developed mathematical models to characterize the electrovibration-induced finger friction and the finger vibrations it causes. Finally, we analyzed how the models’ parameters change across different sliding speeds and applied force settings.
Our results demonstrate that the cutoff frequency of electrovibration-induced friction response increased with sliding speed (Figure 6(b)). We attribute this increase to changes in skin mechanics with sliding speed. In our measurements, both the finger moving mass and stiffness increased with sliding speed and consequently influenced the natural frequency of the finger dynamics, $\omega_{finger}=\sqrt{{k}/{m}}$; check Figure 7(a) and 7(c). While higher stiffness raises the resonance frequency, the increased moving mass lowers it, potentially reducing the cutoff frequency. However, our GLMM analysis revealed that the variation in the stiffness was stronger than the change in mass, resulting in an overall increase in the cutoff frequency. For every 1 mm/s increase in sliding speed, the predicted finger stiffness increased by 4.036 N/m while the finger moving mass increased by $3.223\times 10^{-5}$ kg. Dividing these changes by their measured values at lower speeds, they correspond to increases of 0.82% in mass and 1.11% in stiffness. Additionally, Figure 8) indicates that the correlation coefficient of finger stiffness with cutoff frequency is 30% higher than the correlation with the finger moving mass. These observations provide further evidence that the finger acts less elastic with sliding speed, causing an increase in the cutoff frequency.
The increase in finger stiffness with sliding speed can be attributed to additional shear forces generated during hand motion [37, 38]. When an alternating voltage is applied to the electrovibration touchscreen, the finger is repeatedly attracted to and released from the surface. As the voltage approaches zero, the attraction force diminishes, allowing the shear strain to pull the finger back—a phenomenon that intensifies with increasing finger speed [9]. As the shear strain amplitude increases, finger stiffness increases, decreasing its elasticity [39, 40, 41]. As a result, less energy is dissipated at higher frequencies, as more vibration energy is transmitted into the bulk of the finger [42]. Additionally, amplitude-modulated input further facilitates the vibration energy transmission [43] to the finger bulk, collectively contributing to an increase in the cutoff frequency of the finger’s electrovibration friction response.
Unlike the cutoff frequency, our results (Figure 6(a)) showed that the gain of the electrovibration-induced finger friction was not affected by the sliding speed. We explain this invariance with the observed simultaneous variation in finger moving mass and damping. Our correlation results revealed that the gain decreased with increasing finger moving mass, but increased with increasing finger damping (see Figure 8(a) and Figure 8(c)). A higher finger moving mass implies a lower finger vibration amplitude at low frequencies, which is consistent with our results. Although the finger moving mass increased with speed, the increase is $3.223\times 10^{-5}$ kg per mm/s speed, which is small enough that it does not drastically affect the gain. Similar to the finger moving mass, an increase in finger damping reduces the finger vibration intensity at the resonance. However, our results showed that, like the gain, finger viscosity also remained constant across the sliding speed.
Our results showed that applied force did not have a significant effect on the electrovibration-induced friction and on the finger vibration caused by it (see Figures 6(a)6(b)7(a)7(b), and  7(c)). Interestingly, previous studies have reported an increase in generated friction force due to electrovibration with applied force [27]. Similarly, earlier studies [15, 44, 16] observed a rise in the finger’s mechanical parameters, such as mass, damping, and stiffness with applied force. This contradiction between our results and earlier studies could be caused by the low force ranges for the active applied forces ($0.2\sim 0.6$ N) in our study  [45]. Vardar and Kuchenbecker [13] also found no significant change in the relationship between the input voltage and electrovibration force with varying applied force. They argued that normal forces at 0.5 and 1.0 N still represent sufficiently light and moderate touches not to cause any change in the skin mechanics, which is consistent with our results.
Our findings revealed significant variability in finger skin mechanics among participants, likely due to differences in finger and skin biomechanics, skin moisture, and finger size [46, 47, 48, 49, 50]. Interestingly, the finger electrovibration response parameters ($K$ and $\omega_{o}$) did not show a significant between-participant effect. Since the finger’s mechanical response reflects dynamics across the entire fingertip (from fingerpad to fingernail), while the friction response focuses only on contact changes, it might be concluded that the contact variations were not substantial enough to be detected in the analysis. Also, the skin mechanics and finger-electrovibration friction models used in this study likely oversimplify the complex finger-electrovibration interaction, meaning random effects may not influence them equally.
Our results indicate that at low applied forces ($0.2$ N to $0.6$ N), the spectrum for the electrovibration-induced finger friction remains unchanged. However, as sliding speed increases, the higher-frequency signals of the friction spectrum experience less attenuation, making these frequencies more pronounced. If this speed-dependent behavior is not accounted for in electrovibration-based texture rendering, users may perceive virtual textures differently from the intended ones (e.g., finger friction between real textures or a designed virtual slider in a car cockpit). To ensure a more realistic and robust experience, texture rendering algorithms should dynamically adjust any high-frequency signal based on the sliding speed. With these insights, we proposed an empirically derived model (Equation (7)) that can be used to predict the evolution of finger friction generated by electrovibration based on finger sliding speed for the given input signal magnitude. Using this knowledge, researchers can incorporate the corresponding inverse electrovibration-induced finger friction filter in texture rendering algorithms that continuously minimize the error between the intended texture spectrum and the generated friction spectrum, dynamically adjusting for the input voltage to output friction variations for the finger electrovibration exploration behavior. In the future, we aim to use the proposed finger friction filter and develop a closed-loop texture rendering algorithm that adapts to user exploration, enabling a more natural and free texture exploration experience. Moreover, we aim to evaluate the practical electromechanical response to test whether participants can perceptually differentiate it from user- and condition-specific compensations.
Despite our careful investigation, our work still has limitations. We asked participants to perform a constrained active finger exploration by following the cues in the experimental GUI. This task was quite difficult for some participants, and we had to remove recorded data from seven participants. The participants also found maintaining a consistent contact angle and contact area challenging. Also, the participants performed the experiments in multiple sessions, which could have added to the variability in the model parameters. The potential variations in the contact parameters may have influenced the air gap between the finger and the screen and, consequently, the measured forces. This variability could also explain the lack of a significant trend between gain and sliding speed. It can be reduced when the participant’s finger exploration is controlled with motorized linear stages in the experimental setup.
5 Conclusions
Our results show that finger stiffness increases more than moving mass as sliding speed increases, leading to a rise in the cutoff frequency of the electrovibration-induced friction response. We also observed that low and moderate applied forces do not influence skin mechanics or finger electrovibration response. Building on these insights, we developed a practical, speed-dependent electrovibration-induced friction model that dynamically adjusts the input voltage to achieve the desired output friction. This model enables researchers to enhance the realism of texture rendering algorithms and ensure consistent haptic feedback for users by minimizing the discrepancy between the input intended texture spectrum and the output electrovibration-induced friction spectrum.
6 Data availability
The data and code used in this paper will be publicly available upon acceptance.
7 Acknowledgements
The authors thank Telesilla Bristogianni for helping us with cutting the 3M touchscreen. The authors also thank Michaël Wiertlewski for lending the force sensor for our experimental setup and Celal Umut Kenanoğlu for his feedback on the results and discussion sections.
8 Declaration of competing interest
The authors declare they have no known competing financial interests or personal relationships that could have influenced this work.
9 Funding statement
This work was partly funded by the Dutch Research Council (NWO) with project number 19153.
10 CRediT authorship contribution statement
Jagan K. Balasubramanian: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Writing - Original Draft, and Visualization.
Daan M. Pool: Methodology, Formal Analysis, Writing - Review & Editing.
Yasemin Vardar: Conceptualization, Methodology, Formal Analysis, Writing - Original Draft, Writing - Review & Editing, Supervision, and Funding Acquisition.
Appendix A Electrovibration force generation
A.1 Voltage doubling
This section explains how the input message signal, $V_{m}(t)$, causes electrostatic force $F_{e}(t)$ at twice the message signal’s frequency. For simplicity, let us assume the input is a pure sine wave with frequency $f_{m}$.
$V_{m}(t)=V_{m}\cos(2\pi f_{m}t)$ (8)
From Vardar et al [13], we know $F_{e}\propto kV_{m}^{2}$, where $k$ is a constant. Substituting this relation with the previous equation, we get
$F_{e}(t)\propto k(V_{m}\cos(2\pi f_{m}t))^{2}$ (9)
$F_{e}(t)\propto kV_{m}^{2}\frac{1+\cos(2\times 2\pi f_{m}t)}{2}$ (10)
The equation above shows that the generated electrostatic force has twice the frequency of the input signal.
A.2 Generated friction force through amplitude modulated input signal in electrovibration
This section explains how an electrovibration stimulus generates finger friction through an amplitude-modulated input signal. We first take the square root of the input voltage signal, $V_{m}(t)$, to remove the nonlinearity. Additionally, we DC shift the input voltage signal with $\min|V_{m}(t)|$ to ensure the amplitude is real. The envelope for the amplitude-modulated signal is $V_{env}(t)$ is given as:
$V_{env}(t)=\sqrt{V_{m}(t)+\min|V_{m}(t)|}$ (11)
For amplitude modulation, we multiply the above equation by a high-frequency carrier wave, $f_{c}$, usually beyond 5 kHz. The DC shift, $\min|V_{m}(t)|$, causes the amplitude modulation to be a double-sideband full carrier. The peak amplitude of $V_{m}(t)$ and carrier is 1V, resulting in a modulation index of 1. The modulation index should be $<=1$ to prevent signal distortion due to over-modulation. We chose the modulation index as unity to ensure the input message signal’s power is higher. The carrier signal does not contain useful information for texture perception since the electrovibration friction intensity depends on the input voltage signal’s, $V_{i}(t)$, power.
$V_{i}(t)=\sqrt{V_{m}(t)+\min|V_{m}(t)|}\cos(2\pi f_{c}t)$ (12)
The generated electrovibration force and the input voltage are related through the following equation.
$F_{e}=kV_{i}(t)^{2},$ (13)
where, $F_{e}$ is the electrovibration force, and $k$ is a proportionality constant .
$F_{e}=k\sqrt{V_{m}(t)+\min|V_{m}(t)|}\cos(2\pi f_{c}t))^{2}$ (14)
$F_{e}=k(V_{m}(t)+\min|V_{m}(t)|)\frac{1+\cos(2\pi f_{c}t)}{2}$ (15)
$F_{e}=\frac{k}{2}(V_{m}(t)+\min|V_{m}(t)|+(V_{m}(t)+\min|V(t)|)\cos(2\pi f_{c}t))$ (16)
$F_{e}=\frac{k}{2}(V_{m}(t)+\min|V_{m}(t)|+V_{m}(t)\cos(2\pi f_{c}t)+\min|V_{m}(t)|\cos(2\pi f_{c}t))$ (17)
Substituting $V_{m}(t)=V_{m}cos(2\pi f_{m}t)$ in the above equation.
$F_{e}=\frac{k}{2}(V_{m}\cos(2\pi f_{m}t)+\min|V_{m}(t)|+V_{m}\cos(2\pi f_{m}t)\cos(2\pi f_{c}t)+$ (18)
$\min|V_{m}(t)|\cos(2\pi f_{c}t))$
$F_{e}=\frac{k}{2}(V_{m}\cos(2\pi f_{m}t)+\min|V_{m}(t)|+V_{m}(\cos(2\pi f_{c}+f_{m}t)+$ (19)
$\cos(2\pi f_{c}-f_{m}t))+\min|V(t)|\cos(2\pi f_{c}t))$
Since the finger acts as a low-pass filter with a cut-off at 1 kHz, the higher-frequency components are attenuated:
[…]
The resulting friction force, $f$, from the electrostatic force, $F_{e}$ and the finger normal force, $F_{n}$, is expressed as:
$f=\mu(F_{e}+F_{n})=\mu(\frac{k}{2}(A_{m}\cos(2\pi f_{m}t)+\min|V_{m}(t)|)+F_{n}),$ (21)
Where $\mu$ is the friction coefficient.
Appendix B Appendix B
B.1 Setup response
We wanted to ensure that the experimental setup has a flat response till 1000 Hz as the human tactile range is from 1 Hz to 1000 Hz [30]. We used an impact hammer to excite the system in the normal and lateral directions to measure the response of the experimental setup. The response along the normal direction was found by deriving the transfer function of the setup with vibration as velocity signal, $v_{snd}$, from the impact hammer as output, and the input was recorded from the force sensor, $F_{snd}$. We used $v_{snd}$ as input and $F_{snd}$ as output to make the transfer function causal. In the normal direction, we observed that the setup had a flat response till 500 Hz and a resonance at 1.4 kHz. We fit a second-order model to the setup response in the normal direction, $H(s)_{snd}$ defined as
Figure 9: Frequency response of the experimental setup measured in the normal direction. The setup was excited with an impact hammer. The output response along the normal direction was measured using a force sensor.
Figure 10: Response of the setup measured in the lateral direction. The output response was measured along the force sensor’s lateral axis.
$\frac{v_{snd}(s)}{F_{snd}(s)}=H_{snd}(s)=\frac{sK_{snd}\omega_{n,snd}^{2}}{s^{2}+2\zeta_{snd}\omega_{n,snd}s+\omega_{n,snd}^{2}},$ (22)
Here, $K_{snd}$ is the gain, $\omega_{n,snd}$ is the natural frequency, and $\zeta_{snd}$ is the damping coefficient along the normal direction. By fitting the model to the raw data, we determined the parameters as $K_{snd}=0.58$, $\omega_{n,snd}=1454$ Hz, $\zeta_{snd}=0.011$. We believe the double-sided tape attenuated the measured force, which we accounted for with $K_{snd}$. When measuring the force sensor output without the touchscreen and tape, the magnitude was $10^{0}$.
In the lateral direction, we measured two resonances at 866 Hz and 1740 Hz. We think that the use of double-sided mounting tape caused the first resonance. When measuring the force sensor’s response without the touchscreen and the tape, we detected a single resonance at 2.4 kHz, consistent with the sensor’s datasheet. Hence, we modeled the setup as a fourth-order system, incorporating the mounting effects, as described by the following equation:
$\frac{v_{sld}(s)}{F_{sld}(s)}=H_{sld}(s)=\frac{s\omega_{n,sld}^{2}}{s^{2}+2\zeta_{sld}\omega_{n,sld}s+\omega_{n,sld}^{2}}\cdot\frac{b_{3}s^{2}+b_{2}s+b_{1}}{a_{3}s^{2}+a_{2}s+a_{1}}$ (23)
Here, $v_{sld}(s)$ and $F_{sld}(s)$ are the Fourier transforms of the velocity signal from the impact hammer and lateral force signal from the force sensor, $\zeta_{sld}$ represents the damping coefficient, and $\omega_{n,sld}$ denotes the natural frequency in the lateral direction. The mounting effects are modeled by the transfer function $\frac{b_{3}s^{2}+b_{2}s+b_{1}}{a_{3}s^{2}+a_{2}s+a_{1}}$. Figure 10 illustrates the system’s lateral response. The measured values were $\omega_{n,sld}=1714$ Hz, $\zeta_{sld}=0.06$, $b_{3}=1$, $b_{2}=92$, $b_{1}=(2\pi\cdot 922)^{2}$, $a_{3}=1$, $a_{2}=0.0001$, and $a_{1}=(2\pi\cdot 856)^{2}$.
B.2 Removing setup response
We noticed that the input message signal to output friction raw response and the setup’s response had a first-order behavior, as shown by the sweep markers in Figure 4. The equation gives the response:
$\frac{f(s)}{V_{m}(s)}=H_{f-e}(s)\cdot H_{sld}(s)$ (24)
Here, $f(s)$ and $V_{m}(s)$ are the Fourier transforms of the friction signal from the force sensor and the input voltage envelope. $H_{f-e}(s)$ is the transfer function of finger electrovibration interaction and $H_{sld}$ is the transfer function of the setup’s response. After removing the setup’s response, $H_{sld}$, from the Equation (24), we get a first-order response for $H_{f-e}$ which is modeled as:
$\frac{f(s)}{V_{m}(s)}=H_{f-e}(s)=\frac{K\omega_{o}}{s+\omega_{o}}$ (25)
Here $K$ is the system gain, and $\omega_{o}$ is the cutoff frequency.